OpenStudy (anonymous):

Say I bought a TV for $289.99. I have a deferment plan. The minimum total purchase to get the plan is $149 and up. I picked the 6 month financing. My APR is 25.24%. And the minimum payment requirment is $25.... So this is my scenerio. can someone help me answer the following questions? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what questions are there?

OpenStudy (anonymous):

yay! ok hold on a sec :)

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

ok one sec

OpenStudy (anonymous):

ok :))

jimthompson5910 (jim_thompson5910):

I'm curious if you've learned the remaining balance of a loan formula at all?

OpenStudy (anonymous):

ummm i think so one sec

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

heres a part of the lessons i snipped out. im pretty sure this is it.

jimthompson5910 (jim_thompson5910):

are there any more formulas? I see one that will help with the monthly payment, but I don't see a remaining balance formula

OpenStudy (anonymous):

let me double check

jimthompson5910 (jim_thompson5910):

btw the formula at the very top will be useful for part 2 (since you don't pay down any of the debt, which makes calculating the balance much easier)

OpenStudy (anonymous):

no there are no other formulas

OpenStudy (anonymous):

here are snapshots of the lesson summary to maybe help?

jimthompson5910 (jim_thompson5910):

sorry got distracted, one sec

jimthompson5910 (jim_thompson5910):

well I have a remaining balance formula, but it's a bit complicated and I'm not sure you've learned it yet

OpenStudy (anonymous):

well lets try it!

jimthompson5910 (jim_thompson5910):

ok the formula is B = L*(1+r/n)^k - (P*n/r)*( (1+r/n)^k - 1 ) Where, B = Remaining balance at month k (payment is made on month k) L = Loan amount P = Payment per period r = Annual interest rate n = Compounding frequency k = Point in time where you want to find the remaining balance (time in months)

OpenStudy (anonymous):

no i havnt learned that but i can see what they all mean

jimthompson5910 (jim_thompson5910):

In this case, B = Unknown (we're solving for this) L = 289.99 P = 25 r = 25.24 n = 12 (we're compounding it monthly) k = 6 (you want to find the remaining balance at month 6)

jimthompson5910 (jim_thompson5910):

Plug all these values in to get B = L*(1+r/n)^k - (P*n/r)*( (1+r/n)^k - 1 ) B = 289.99*(1+0.2524/12)^6 - (25*12/0.2524)*( (1+0.2524/12)^6 - 1 ) B = 170.453723678253 B = 170.45

OpenStudy (anonymous):

this is the remaining balance?

jimthompson5910 (jim_thompson5910):

So what this means is that if you buy a TV that costs 289.99 and you make 6 monthly payments during the deferment period, this will mean that you would still owe $170.45 after the deferment period is over

jimthompson5910 (jim_thompson5910):

6 monthly payments of $25*

OpenStudy (anonymous):

this is number 1 right?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thats easier to understand then i expected actually! haha lets move on to number two

jimthompson5910 (jim_thompson5910):

ok now we use the monthly payment formula since it says that we don't have to show work, we can use a monthly calculator instead which method do you want to use?

OpenStudy (anonymous):

we can just use the calculator:) its much easier

jimthompson5910 (jim_thompson5910):

ok great, go here then http://www.bankrate.com/calculators/mortgages/loan-calculator.aspx and type in 170.45 into the loan amount box, 5 in the years box, and 25.24 in the rate box and hit calculate tell me what you get

OpenStudy (anonymous):

ok one sec :)

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

i got 5.03?

jimthompson5910 (jim_thompson5910):

me too

jimthompson5910 (jim_thompson5910):

that's the monthly payment and the answer for #2

jimthompson5910 (jim_thompson5910):

for #3, during the deferment period you paid 25*6 = 150 dollars then after the deferment period, you paid 60*5.03 = 301.8 dollars (since 5 years = 60 months) in total you paid 150 + 301.8 = 451.80 dollars the total cost is therefore 451.80 dollars

jimthompson5910 (jim_thompson5910):

#4 should be pretty straight forward

OpenStudy (anonymous):

so 281.35?

jimthompson5910 (jim_thompson5910):

no you subtract 451.80 - 289.99

jimthompson5910 (jim_thompson5910):

this is to find out how much extra you pay over the long run (if you finance instead of just pay all of the $289.99 at once)

OpenStudy (anonymous):

161.81? :)

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

for the next little section i know im going to have some trouble because thiese questions dont include the minimum payment...

jimthompson5910 (jim_thompson5910):

ok for #1 in the next section, you are going to use the formula that is B = P(1+i)^(nt)

jimthompson5910 (jim_thompson5910):

In this case P = 289.99 i = 0.2524/12 n = 12 t = 6/12 = 0.5

OpenStudy (anonymous):

i didnt get 0.5??

jimthompson5910 (jim_thompson5910):

t is time in years so 6 months = 6/12 = 0.5 years

jimthompson5910 (jim_thompson5910):

that's why t = 0.5

OpenStudy (anonymous):

oh i miscalculated it. now i got it:)

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (anonymous):

ok what do we do for number 2?

jimthompson5910 (jim_thompson5910):

what did you get for #1

OpenStudy (anonymous):

.05?

jimthompson5910 (jim_thompson5910):

no that's not the answer to #1

jimthompson5910 (jim_thompson5910):

use the formula I wrote above

OpenStudy (anonymous):

170.45

jimthompson5910 (jim_thompson5910):

no that was the answer to #1 from part 1

OpenStudy (anonymous):

oh i thought thats what you meant

jimthompson5910 (jim_thompson5910):

no i mean #1 from part 2

OpenStudy (anonymous):

i dont know? i dont know how to find the mpnthly pay ment....

jimthompson5910 (jim_thompson5910):

plug these values P = 289.99 i = 0.2524/12 n = 12 t = 6/12 = 0.5 into the formula below B = P(1+i)^(nt)

OpenStudy (anonymous):

wouldnt that give me the same exact answer since we already plugged these numbers into this formula?

jimthompson5910 (jim_thompson5910):

no it's much higher than 170.45

OpenStudy (anonymous):

ok one sec.

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

so B= 328.566?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

oh wait, hold on

OpenStudy (anonymous):

what? i dont understnad i plugged them all and and solved it...

jimthompson5910 (jim_thompson5910):

sorry made a typo, I'm getting 289.99(1+0.2524/12)^(12*0.5) = 328.565943133299 so you are correct

jimthompson5910 (jim_thompson5910):

so the answer to #1 in part 2 is 328.57

OpenStudy (anonymous):

oh whew! and woops i thought this was the answer to number 2 in part 2 lol

jimthompson5910 (jim_thompson5910):

no you use that calculator you used from part 1

jimthompson5910 (jim_thompson5910):

to answer #2 in part 2

OpenStudy (anonymous):

oh! let me try this one on my own hold on just a sec :)

OpenStudy (anonymous):

oh gosh not even close. i dont know what to plug in

jimthompson5910 (jim_thompson5910):

328.57 in the loan amount box 5 in the years box 25.24 in the rate box

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

$9.69

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

yay!

jimthompson5910 (jim_thompson5910):

to answer #3, you need to multiply the number of months (60) by the monthly payment ($9.69) to get the total cost

OpenStudy (anonymous):

581.4?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

thats the answer to number #3? it seems like we did more work in part 1 number #3? rather than this onw?

jimthompson5910 (jim_thompson5910):

well that's because back in part 1, there's the amount you add on during the deferment period

jimthompson5910 (jim_thompson5910):

but you don't have to worry about that in this part

OpenStudy (anonymous):

oh gotcha! that makes sense:)

OpenStudy (anonymous):

now number 5 :)

jimthompson5910 (jim_thompson5910):

what did you get for #4

OpenStudy (anonymous):

oh my gosh sorry ahah im getting ahead of myself. i meant number 4

jimthompson5910 (jim_thompson5910):

#4 is pretty straight forward, tell me what you get

OpenStudy (anonymous):

38.58!

jimthompson5910 (jim_thompson5910):

you are subtracting the original cost from the total cost 581.4-289.99 = ??

OpenStudy (anonymous):

291.41